# Task Preference Optimization: Improving Multimodal Large Language Models with Vision Task Alignment

### v1

@@ 2 Training and Data Details
Table 17: Training Datasets. The temporal grounding includes two subtasks: moment retrieval and highlight detection.
Table 16 and 17 lists the detailed training configurations and data of VideoChat-TPO in different stages. In each stage, the model is parametrized from the weights from the previous stage and continues training.
@@
In stage 2, the region head and token are trained with a learning rate of 2e-5 using a cosine learning rate scheduler. We use a two-layer MLP with a MSE loss as the region head to train from scratch. For training data, we use AS-V2 [85], Visual Genome [37], RefCOCO [103], RefCOCOg [103], RefCOCO+ [103] for one epoch with a batch size of 128.
@@
For the mask head, we use the pre-trained SAM2 [69] model, replacing the prompt encoder of SAM2 with a single MLP layer (mask adapter). During learning, only the mask token and adapter are trained with a learning rate of 2e-5. We use MeViS [18] , SAMv2 [69] for three epochs in this stage with a batch size of 128. We supplement the training data by adding a expanded ASv2 [85] (we convert images into videos) to this stage.
@@
1. Where is $\langle expr\rangle$? 2. Can you find $\langle expr\rangle$? 3. Can you detect $\langle expr\rangle$? 4. Can you locate $\langle expr\rangle$? 5. Please find $\langle expr\rangle$. 6. Please detect $\langle expr\rangle$? 7. Please locate $\langle expr\rangle$? 8. Find $\langle expr\rangle$. 9. Detect $\langle expr\rangle$? 10. Locate $\langle expr\rangle$?
@@
1. Please give the motion path of $\langle obj\rangle$ in the video over time.
2. Show the tracking trajectory of $\langle obj\rangle$’s movement through the scene in the video.
3. Please generate a motion path of $\langle obj\rangle$’s movement in the video, highlighting its tracking trajectory.
4. Show the tracking trajectory of $\langle obj\rangle$.
5. Generate $\langle obj\rangle$’s tracking trajectory.
6. Visualize the tracking trajectory of $\langle obj\rangle$ in the video.
7. Please generate a visual representation of $\langle obj\rangle$’s movement in the video, highlighting its tracking trajectory.
@@
To support the proper invocation of task-specific decoders, we construct a series of instruction templates for different tasks and use them as instruction tuning data for MLLM. We comprehensively list all the instruction templates below, in Table 21, 21, 21, and 21.
@@ 3 Qualitative Results
In Figure 4, we show the spatial grounding visualizations. VideoChat-TPO finds the target object from the description of natural language. It accurately locates the target among multiple similar objects, even if the target is occluded or in the background area.
@@
Some tracking results are shown in Figure 6. Users need to locate the target (e.g. bounding box coordinates) in the first frame of the input video. The visualizations show that when the target object is partially occluded in the video, it can still be tracked. Even if the target object is out of the camera’s view, VideoChat-TPO can still follow it when it appears in subsequent frames.
@@
The moment retrieval and highlight detection are illustrated in Figure 7. VideoChat-TPO can target events based on the user’s questions, and perform moment retrieval and highlight detection on the target events.
@@
The multimodal video understanding visualizations are shown in Figure 8. VideoChat-TPO achieves decent results in fine-grained action description, spatial description, and video captioning.
@@
Figure 7: Qualitative Results of Moment Retrieval and Highlight Detection. The orange curve represents the saliency score, the blue interval represents the time interval predicted by the model, and the green interval represents the ground truth.

### v2

@@ 2 Training and Data Details
Table 18: Datasets Used at Three Training Stages. The temporal grounding task includes two subtasks: moment retrieval and highlight detection.
Table 17 and 18 lists the detailed training configurations and data of VideoChat-TPO in different stages. In each stage, the model is parametrized from the weights from the previous stage and continues training.
@@
In stage 2, the region head and token are trained with a learning rate of 2e-5 using a cosine learning rate scheduler. We use a two-layer MLP as region head to train from scratch and we use MSE loss for region head training. For training data, we use AS-V2 [83], Visual Genome [36], RefCOCO [98], RefCOCOg [98], RefCOCO+ [98] for one epoch with a total batch size of 128 to train region head and token.
@@
For the mask head, we use the pre-trained SAM2 [68] model, replacing the prompt encoder of SAM2 with a single MLP layer called the mask adapter. During training, the mask token and adapter are trained with a learning rate of 2e-5, and the rest of SAM2 is frozen. We use MeViS [17] , SAMv2 [68] for three epochs in this stage with a total batch size of 128. We supplement the training data by expanding the ASv2 [83] image dataset into videos and adding it to this stage.
@@
TPO Additional Training and Inference Cost.
Model GPU Stage1 Stage2 Stage3
VideoChat-TPO 64 0.5h 11h 52h
        textualized task data 64 0.5h 10h 50h
        only conversation data 64 - - 42h
Table 19: Training Cost of Three Stages on VideoChat. Textualized task data means converting task data into conversation form.
From the data perspective, as can be seen from Table 18 in the Appendix, we have very little training data in the first (around 0.15M) and second stages (around 0.7M), most of the data (around 3.5M) is used in third phase of the experiment. Among the data in the third stage, most of it is conversation data for fine-tuning MLLM. Therefore, TPO introduces little new data. Concerning training cost, according Table  19, when using the same amount of data, the training time of our TPO method and the autoregressive method is almost the same, and compared with the version without visual task, the TPO method increases the training cost by about 25%.
The Temporal Head and the Mask Head contains additional encoders. In training phase, the additional encoders are frozen, and we use the features extracted by the encoder for training. In inference phase, the additional encoders are only used when the task head is activated. When only performing conversation tasks, no additional inference cost is incurred.
@@
1. Where is $\langle expr\rangle$?
2. Can you find $\langle expr\rangle$?
3. Can you detect $\langle expr\rangle$?
4. Can you locate $\langle expr\rangle$?
5. Please find $\langle expr\rangle$.
6. Please detect $\langle expr\rangle$?
7. Please locate $\langle expr\rangle$?
8. Find $\langle expr\rangle$.
9. Detect $\langle expr\rangle$?
10. Locate $\langle expr\rangle$?
@@
1. Please give the motion path of $\langle obj\rangle$ in the video over time. 2. Show the tracking trajectory of $\langle obj\rangle$’s movement through the scene in the video. 3. Please generate a motion path of $\langle obj\rangle$’s movement in the video, highlighting its tracking trajectory. 4. Show the tracking trajectory of $\langle obj\rangle$. 5. Generate $\langle obj\rangle$’s tracking trajectory. 6. Visualize the tracking trajectory of $\langle obj\rangle$ in the video. 7. Please generate a visual representation of $\langle obj\rangle$’s movement in the video, highlighting its tracking trajectory.
@@
To support the proper invocation of task-specific decoders, we construct a series of instruction templates for different tasks and use them as instruction tuning data for MLLM. We comprehensively list all the instruction templates below, in Table 23, 23, 23, and 23.
@@ 3 Qualitative Results
In Figure 4, we show the spatial grounding visualizations. VideoChat-TPO can infer the target object from the description of natural language and locate it. Our VideoChat-TPO can accurately locate the target among multiple similar objects. Even if the target object is occluded or in the background area, it can still be accurately located.
@@
The tracking visualizations are shown in Figure 6. The user needs to include the bounding box coordinate information of the first frame of the tracked target in the video in the input. The visualizations show that when the target object is partially occluded in the video, it can still be tracked. Even if the target object is out of the camera’s view, our VideoChat-TPO can still track it when it appears in subsequent frames.
@@
The visualizations of the moment retrieval and highlight detection are given in Figure 7. Our VideoChat-TPO can infer the results and target events based on the user’s questions, and perform moment retrieval and highlight detection on the target events.
@@
The multimodal video understanding visualizations are shown in Figure 8. Our VideoChat-TPO achieve decent results in fine-grained action description, spatial description, and video captioning.
@@
Figure 7: Qualitative Results of Moment Retrieval and Highlight Detection. The orange curve represents saliency score, the blue interval represents the time interval predicted by the model, and the green interval represents the ground truth.